# Kummer-Artin-Schreier-Witt Theory

### v2

@@ 3.2 Construct the Kummer Extension
    $G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\llbracket X_{0},\ldots,X_{i}\rrbracket$ satisfies the Tate condition,
@@
    $\nu(G_{i+1}^{\prime})>\frac{p}{p-1}$ and $G_{i+1}^{\prime}\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\llbracket X_{0},\ldots,X_{i}\rrbracket$ satisfies the Tate condition, and
@@
Let us first consider the case $i=1$. Then we can simply pick $u_{0}=1+p\lambda X_{0}$ or $G_{1}=1$. By Proposition 3.2, the Witt vector associated with the order $p$ Kummer class $u_{0}$ is $(X_{0}^{\prime})$ where
$\begin{split}X_{0}^{\prime}&=\frac{1}{p\lambda}\ln(1+p\lambda X_{0})\\ &=\frac{1}{p\lambda}\left(p\lambda X_{0}-\frac{p^{2}\lambda^{2}X_{0}^{2}}{2}+\cdots+(-1)^{j+1}\frac{p^{j}\lambda^{j}X_{0}^{j}}{j}+\cdots\right)\end{split}$
The valuations of the coefficients are given by
$\nu\left(\frac{p^{j}\lambda^{j}}{p\lambda j}\right)=\begin{cases}j-1+\frac{j-1}{p-1}=\frac{p(j-1)}{p-1},&p\nmid j\\ p^{i}l-1+\frac{p^{i}l-1}{p-1}-i=\frac{p(p^{i}l-1)}{p-1}-i,&j=p^{i}l,\;(p,l)=1\end{cases}$
Hence, it is clear that the coefficient of $X_{0}^{j}$ has valuation greater than 00 when $j>1$, and $X_{0}^{\prime}$ satisfies the Tate condition. This completes the base case.
Now, consider the case $i=s$. By Lemma 2.4, $G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))$, regarded as a power series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0}^{\prime},\ldots,X_{s-2}^{\prime}\rrbracket$, satisfies the Tate condition. Furthermore, by the induction hypothesis, $X_{j}^{\prime}$ satisfies the same condition for $j=0,\ldots,s-2$ as a series over $\mathbb{Z}_{(p)}[\zeta_{p^{j+1}}]\llbracket X_{0},\ldots,X_{j}\rrbracket$. Therefore, $G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))$ satisfies the condition as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\ldots,X_{s-2}\rrbracket$, proving 1.
Next, we obtain $G_{s}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}][X_{0},\ldots,X_{s-2}]$ by truncating $G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))$ at terms with valuations at most $\frac{p}{p-1}$. Assertion 2 is immediately verified. Assertion 3 can be derived from the computation of $X_{s-1}^{\prime}$ in the proof of Proposition 3.3. Recall that
$X_{s-1}^{\prime}$ $=\frac{1}{p\lambda}\ln\left(1+\frac{-G^{\prime}_{s}+p\lambda X_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right)=:\frac{1}{p\lambda}\ln(1+U).$
It is clear that both the numerator and the denominator of $U$ satisfy the condition, so $U$ itself does as well. The claim that $\ln(1+U)$ verifies the condition follows immediately. ∎
@@ 3.3 Defining the étale Cover
Define $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$ and set $\pi:=\zeta_{p^{s}}-1$. Let $\mathcal{V}_{s}$ be an open subset of the projective space $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ given by
@@
where each $G_{i}$ is a polynomial constructed in Proposition 3.5. Let $\mathcal{W}_{s}$ be the integral closure of $\mathcal{V}_{s}$ over the extension of $K(X_{0},\ldots,X_{s-1})$ defined by
$Z_{s-1}^{p^{s}}=(1+p\lambda X_{0})\prod_{i=1}^{s-1}(G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i})^{p^{i}}.$ (19)
@@
The special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is birational to the Artin-Schreier-Witt extension of $\operatorname{Proj}k[x_{0},\ldots,x_{s-1},y]$, where $x_{i}$ (resp. $y$) is the reduction of $X_{i}$ (resp. $Y$) modulo $\pi$, given by the Witt vector $(x_{0},\ldots,x_{s-1})$.
@@
Applying Lemma 2.6 to $K:=\mathbb{Q}[\zeta_{p^{s}}]\mathopen{(\!(}X_{0},\ldots,X_{s-2}\mathclose{)\!)}$, which is discretely valued with respect to the Gauss valuation, and setting $z=X_{s-1}$, we obtain a Henselian ring $\mathcal{R}^{\operatorname{int}}$ and a Henselian field $\mathcal{R}^{\text{bd}}$ containing $\mathbb{Q}[\zeta_{p^{s}}](X_{0},\ldots,X_{s-1})$ with residue field $\kappa=\mathbb{F}_{p}\mathopen{(\!(}x_{0},\ldots,x_{s-1}\mathclose{)\!)}$.
It follows from the proof of Proposition 3.3 that $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})$ lies in $W_{s}(\mathcal{R}^{\operatorname{int}})$. Therefore, by Theorem 2.7, the extension of $\mathcal{R}^{\text{bd}}$ given by (19) corresponds to the Witt vector $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})$, which is a lift of $(x_{0},\ldots,x_{s-1})\in W_{s}(\kappa)$. This completes the proof. ∎
The following result follows directly from Kummer theory.
@@
Specifically, the cover of $\mathbb{P}^{s}_{R}$ defined by (19) is branched precisely along the hypersurfaces given by $G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i}=0$ for $10\leq i\leq s-$ 1. Consequently, the cover $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is unramified.
4 Compute the $\mathcal{W}_{s}$
@@ 4.1 The Case $s=1$
Consider the change of variables $Z_{0}=1+\lambda Y_{0}$. This substitution transforms (20) into
@@
[13, Section 2.4], the equation above reduces modulo $\lambda$ to the Artin-Schreier equation
@@
To generalize this change of variables for the case $s>1$, we follow the strategy of associating a Witt vector as in the previous section. Recall that the right-hand side of (20), due to Proposition 3.3, is associated with the Witt vector $(X_{0}^{\prime})$, where $X_{0}^{\prime}\in\mathbb{Z}_{(p)}[\zeta_{p}]\llbracket X_{0}\rrbracket$ is congruent to $X_{0}$ modulo $\lambda$. We can rewrite the defining equation as
@@ 4.2 Inductive Step
to better understand the structure of $\mathcal{W}_{s}$ and to derive an Artin-Schreier-type equation similar to (21). As in Section 4.1, beginning with the system of $\mathbb{Z}/p$-Kummer extensions
$Z_{i-1}^{p}=Z_{i-2}(G_{i}(X_{0},\ldots,X_{i-2})+p\lambda X_{i-2}),$
we seek a Witt vector $(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$ such that, for $1\leq i\leq s$,
@@
The desired formula for $Y_{s-1}^{\prime}$ follows immediately. ∎
@@ 4.2.1 Approximation of the Power Series $E_{s,p}$
All terms of $E_{i,p}((X_{0},\ldots,X_{i-1}))$, aside from the constant term $1$, have positive valuations that are sharply bounded below by $\frac{1}{p^{i}(p-1)}$.
@@
We begin by computing the valuations for the case $i=1$. Then
$\begin{split}E_{1,p}((X_{0}))&=\exp(\lambda X_{0})\\ &=1+\sum_{i\geq 1}\frac{\lambda^{i}X_{0}^{i}}{i!}.\end{split}$
It is well-known that $\nu(n!)=\sum_{j\geq 1}\left\lfloor\frac{n}{p^{j}}\right\rfloor=\frac{n-s_{p}(n)}{p-1}$, where $s_{p}(n)$ denotes the sum of digits of $n$ in base $p$. Therefore, the valuations of the coefficients are given by
$\nu\left(\frac{\lambda^{i}}{i!}\right)=\begin{cases}\frac{i}{p-1}-\frac{i-s_{p}(i)}{p-1}=\frac{s_{p}(i)}{p-1}\geq\frac{1}{p-1},&p\nmid i,\\ \frac{p^{n}}{p-1}-\frac{p^{n}-1}{p-1}=\frac{1}{p-1},&i=p^{n}.\end{cases}$
This proves the base case, so we have $\nu(E_{1,p}((X_{0}))-1)=\frac{1}{p-1}$.
For $i>1$, using the commutativity of $E_{i,p}$, we find
$\begin{split}E_{i,p}(p(X_{0},\ldots,X_{i-1}))&=E_{i,p}((pX_{0},0,\ldots,0))\cdot E_{i,p}((0,\tilde{X}_{1},\ldots,\tilde{X}_{i-1}))\\ &=E_{i,p}((pX_{0},0,\ldots,0))\cdot E_{i-1,p}((\tilde{X}_{1},\ldots,\tilde{X}_{i-1})),\end{split}$
where $\tilde{X}_{j}=X_{j-1}^{p}+pR$. By the induction hypothesis, $\nu(E_{i-1,p}((\tilde{X}_{1},\ldots,\tilde{X}_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}$, which is strictly less than 1 in general, except when $p=2$ and $i=2$. Thus, $\nu(E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}$ for $p>2$ or $p=2$ and $i>2$.
Now, suppose $E_{i,p}((X_{0},\ldots,X_{i-1}))=1+A$, where $A\in R:=\mathbb{Z}_{(p)}[\zeta_{p^{i}}]\llbracket X_{0},\ldots,X_{i-1}\rrbracket$. Then
$E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1=A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j}.$
Since $\nu(E_{i,p}(p(X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-2}(p-1)}<1$ unless $p=2$ and $i=2$, we deduce that $\nu(A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j})=p\nu(A)$. Therefore, we conclude that $\nu(E_{i,p}((X_{0},\ldots,X_{i-1}))-1)=\frac{1}{p^{i-1}(p-1)}$.
Finally, a direct computation confirms that $\nu(E_{2,2}((X_{0},X_{1}))-1)=\frac{1}{2}$, completing the proof. ∎
@@
Suppose $f(X)=1+\sum_{j\geq 0}a_{j}X^{j}\in R\llbracket X\rrbracket$, with $\nu(a_{j})>0$ and the valuations bounded below by a number $B$. Suppose moreover that $f(X)^{p}$ has a finite number of terms whose coefficients have valuations strictly smaller than $p/(p-1)$. Then $f(X)$ itself has only finitely many terms whose coefficients have valuations strictly smaller than $1/(p-1)$.

### v3

@@ 3.2 Constructing the Kummer Extension
    $G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\llbracket X_{0},\ldots,X_{i-1}\rrbracket$ satisfies the Tate condition,
@@
    $\nu(G_{i+1}^{\prime})>\frac{p}{p-1}$ and $G_{i+1}^{\prime}\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\llbracket X_{0},\ldots,X_{i-1}\rrbracket$ satisfies the Tate condition, and
@@
We proceed by induction on $i\geq 0$.
Let’s consider the base case $i=0$. Take $G_{1}=1$, so $u_{0}=1+p\lambda X_{0}$. By Proposition 3.4, the associated Witt coordinate is
$X_{0}^{\prime}=\frac{1}{p\lambda}\ln(1+p\lambda X_{0})=\sum_{j\geq 1}(-1)^{j+1}\frac{(p\lambda)^{j-1}}{j}X_{0}^{j}.$
Writing $j=p^{r}\ell$ with $(p,\ell)=1$ and using $\nu(\lambda)=\frac{1}{p-1}$ and $\nu(j)=r$ (since $\nu(p)=1$), we have
$\nu\!\left(\frac{(p\lambda)^{j-1}}{j}\right)=(j-1)\,\nu(p\lambda)-\nu(j)=(j-1)\left(1+\frac{1}{p-1}\right)-r=\frac{p(j-1)}{p-1}-r.$
In particular, as $j\to\infty$ we have
$\nu\!\left(\frac{(p\lambda)^{j-1}}{j}\right)=\frac{p(j-1)}{p-1}-r\longrightarrow\infty,$
since $r=\nu(j)=\nu_{p}(j)$ grows at most logarithmically in $j$. Hence $X_{0}^{\prime}$ satisfies the Tate condition, establishing the base case.
Assume that $G_{1},\ldots,G_{i}$ have been constructed and that $X_{0}^{\prime},\ldots,X_{i-1}^{\prime}$ satisfy the Tate condition.
By Lemma 2.5, the series $G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))$ satisfies the Tate condition when viewed as a power series in the variables $X_{0}^{\prime},\ldots,X_{i-1}^{\prime}$. Since each $X_{j}^{\prime}$ itself satisfies the Tate condition as a power series in $X_{0},\ldots,X_{j}$ and has zero constant term, we may substitute these series into $G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))$. By Lemma 3.7, the result again satisfies the Tate condition as a series in $X_{0},\ldots,X_{i-1}$. This proves 1.
Define $G_{i+1}(X_{0},\ldots,X_{i-1})$ to be the truncation of $G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))$ obtained by keeping only those terms whose coefficients have valuation at most $\frac{p}{p-1}$. Then $G_{i+1}^{\prime}$ has coefficientwise valuation strictly larger than $\frac{p}{p-1}$, and it still satisfies the Tate condition. This proves 2.
Finally, by Proposition 3.4 we have
$X_{i}^{\prime}=\frac{1}{p\lambda}\ln\left(\frac{G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i}}{G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))}\right)=\frac{1}{p\lambda}\ln(1+U),$
where
$U=\frac{-G_{i+1}^{\prime}+p\lambda X_{i}}{G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))}.$
The numerator and denominator of $U$ satisfy the Tate condition, and the denominator is a unit (with constant term $1$), so its inverse also satisfies the Tate condition; hence $U$ satisfies the Tate condition in $\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\llbracket X_{0},\ldots,X_{i}\rrbracket$.
Now write
$\ln(1+U)=\sum_{m\geq 1}(-1)^{m-1}\frac{U^{m}}{m}.$
Set $\delta:=\nu(U)>0$. Fix $M\in\mathbb{R}$. Choose $m_{0}$ such that
$m\delta-\nu(m)>M\qquad(m\geq m_{0}).$
Then for $m\geq m_{0}$, every coefficient of $U^{m}/m$ has valuation >M>M. For each of the finitely many $1\leq m<m_{0}$, since $U^{m}$ satisfies the Tate condition, there exists $B_{m}$ such that all coefficients of $U^{m}/m$ of total degree $B_{m}$ have valuation >M>M. Let $B=\max_{1\leq m<m_{0}}B_{m}$. Then every coefficient of $\ln(1+U)$ of total degree >B>B has valuation >M>M. Therefore $\ln(1+U)$ satisfies the Tate condition. It follows that $X_{i}^{\prime}$ satisfies the Tate condition as well, proving 3. ∎
@@ 3.3 Defining the Étale Cover
Define $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$ and set $\pi:=\zeta_{p^{s}}-1$. Let $\mathcal{V}_{s}$ be the open subset of the projective space $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ given by
@@
where each $G_{i}$ is a polynomial constructed in Proposition 3.8. Let $\mathcal{W}_{s}$ be the integral closure of $\mathcal{V}_{s}$ in the extension of $K(X_{0},\ldots,X_{s-1})$ defined by
$Z_{s-1}^{p^{s}}=(1+p\lambda X_{0})\prod_{i=1}^{s-1}\big(G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i}\big)^{p^{i}}.$ (21)
@@
The special fiber of $\mathcal{W}_{s}\to\mathcal{V}_{s}$ is birational to the Artin–Schreier–Witt cover of $\operatorname{Proj}k[x_{0},\ldots,x_{s-1},y]$ associated with the Witt vector $(x_{0},\ldots,x_{s-1})$, where $x_{i}$ (resp. $y$) is the reduction of $X_{i}$ (resp. $Y$) modulo $\pi$.
@@
Applying Proposition 3.1 with $K=\mathbb{Q}(\zeta_{p^{s}})$, we obtain a Henselian local field $\mathbb{K}=\mathcal{R}^{\mathrm{bd}}$ with associated ring $\mathcal{R}^{\mathrm{int}}$ and residue field $\kappa=\mathbb{F}_{p}((x_{0},\ldots,x_{s-1}))$.
By construction, the Witt vector $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})$ is obtained from $(X_{0},\ldots,X_{s-1})$ by iterated application of the formal series $\ln(1+\cdot)$ (and multiplication by $(p\lambda)^{-1}$). Hence each $X_{i}^{\prime}$ is a formal power series in $X_{0},\ldots,X_{i}$ with coefficients in $R=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$. By Lemma 3.2, we may view these series as elements of $\mathcal{R}^{\mathrm{int}}$. In particular, $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\in W_{s}(\mathcal{R}^{\mathrm{int}})$.
Therefore, by Theorem 2.8, the Kummer class defined by (21) corresponds to the Witt vector $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})$. Reducing modulo $\pi$ gives $(x_{0},\ldots,x_{s-1})\in W_{s}(\kappa)$, so the special fiber is birational to the Artin–Schreier–Witt cover associated with $(x_{0},\ldots,x_{s-1})$. ∎
@@
The cover $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is generically unramified.
Proof.
By Kummer theory, the extension of function fields defined by (21) is cyclic of degree $p^{s}$ (hence generically Galois with group $\mathbb{Z}/p^{s}$). Its branch locus on $\mathbb{P}^{s}_{R}$ is contained in the complement of the maximal open subset on which the Kummer class on the right-hand side of (21) is a unit. Concretely, it is contained in the union of the divisor at infinity
$Y=0$
(taking $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$) together with the divisors
$G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i}=0\quad(0\leq i\leq s-1).$
By definition of $\mathcal{V}_{s}\subset D_{+}(Y)$, each $G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i}$ is inverted on $\mathcal{V}_{s}$, hence is a unit there; also $\mathcal{V}_{s}\subset D_{+}(Y)$ is disjoint from $Y=0$. Therefore the branch locus is disjoint from $\mathcal{V}_{s}$, so the induced morphism $\mathcal{W}_{s}\to\mathcal{V}_{s}$ is generically unramified. ∎
4 Computing the $\mathcal{W}_{s}$
@@ 4.1 The case $s=1$
Consider the change of variables $Z_{0}=1+\lambda Y_{0}$. This transforms (22) into
@@
[25, Section 2.4], the equation above reduces modulo $\lambda$ to the Artin–Schreier equation
@@
To generalize this change of variables to the case $s>1$, we follow the strategy of associating a Witt vector as in the previous section. Recall that the right-hand side of (22), due to Proposition 3.5, is associated with the Witt vector $(X_{0}^{\prime})$, where $X_{0}^{\prime}\in\mathbb{Z}_{(p)}[\zeta_{p}]\llbracket X_{0}\rrbracket$ is congruent to $X_{0}$ modulo $\lambda$. We can rewrite the defining equation as
@@ 4.2 Inductive step
to better understand the structure of $\mathcal{W}_{s}$ and to derive an Artin–Schreier-type equation similar to (23). As in Section 4.1, beginning with the system of $\mathbb{Z}/p$-Kummer extensions
$Z_{i-1}^{p}=Z_{i-2}\big(G_{i}(X_{0},\ldots,X_{i-2})+p\lambda X_{i-1}\big),$
we seek a Witt vector $(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$; equivalently, we write
$(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})=(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$
for the same Witt vector viewed in the $X$-coordinates on $\mathcal{V}_{s}$. Then, for $1\leq i\leq s$,
@@
The desired formula for $Y_{s-1}^{\prime}$ follows immediately. Replacing $s$ by $i+1$ in the same computation gives the displayed formula for each $Y_{i}^{\prime}$. ∎
@@ 4.2.1 Approximation of the Power Series $E_{s,p}$
All terms of $E_{i,p}((X_{0},\ldots,X_{i-1}))$, aside from the constant term $1$, have positive valuations that are sharply bounded below by $\frac{1}{p^{i-1}(p-1)}$.
@@
We write $\nu(f)$ for the minimum valuation among the coefficients of a (nonzero) power series $f\in\mathbb{Z}_{(p)}[\zeta_{p^{i}}]\llbracket X_{0},\ldots,X_{i-1}\rrbracket$. Equivalently, $\nu(f-1)$ is the smallest valuation among the nonconstant terms of $f$.
Step 1: the case $i=1$. We have
$E_{1,p}((X_{0}))=\exp(\lambda X_{0})=1+\sum_{n\geq 1}\frac{\lambda^{n}}{n!}X_{0}^{n}.$
Using Legendre’s formula $\nu(n!)=\sum_{j\geq 1}\lfloor n/p^{j}\rfloor=\frac{n-s_{p}(n)}{p-1}$ and $\nu(\lambda)=\frac{1}{p-1}$, we obtain
$\nu\!\left(\frac{\lambda^{n}}{n!}\right)=\frac{n}{p-1}-\frac{n-s_{p}(n)}{p-1}=\frac{s_{p}(n)}{p-1}\geq\frac{1}{p-1}.$
Hence every nonconstant coefficient of $E_{1,p}$ has valuation $\geq\frac{1}{p-1}$, and the bound is sharp (e.g. for $n=1$).
Step 2: reduction to the valuation of $E_{i,p}(p\underline{X})-1$. Fix $i>1$ and write $\underline{X}=(X_{0},\ldots,X_{i-1})$. Set
$E_{i,p}(\underline{X})=1+A,\qquad A\in\mathbb{Z}_{(p)}[\zeta_{p^{i}}]\llbracket X_{0},\ldots,X_{i-1}\rrbracket.$
Since $E_{i,p}$ is a homomorphism from the additive group of Witt vectors to units (cf. (15)), we have
$E_{i,p}(p\underline{X})=E_{i,p}(\underline{X})^{p}=(1+A)^{p}.$
Therefore
$E_{i,p}(p\underline{X})-1=A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j}.$
Step 3: compute $\nu\big(E_{i,p}(p\underline{X})-1\big)$. In the Witt ring, one has the standard decomposition
$p\underline{X}=(pX_{0},0,\ldots,0)+(0,\widetilde{X}_{1},\ldots,\widetilde{X}_{i-1}),$
where each $\widetilde{X}_{j}\equiv X_{j-1}^{p}\pmod{p}$ (more precisely, $\widetilde{X}_{j}=X_{j-1}^{p}+p\cdot(\text{a series}))$. Using again multiplicativity of $E_{i,p}$, we get
$E_{i,p}(p\underline{X})$ $=E_{i,p}((pX_{0},0,\ldots,0))\cdot E_{i,p}((0,\widetilde{X}_{1},\ldots,\widetilde{X}_{i-1}))$
$=E_{i,p}((pX_{0},0,\ldots,0))\cdot E_{i-1,p}((\widetilde{X}_{1},\ldots,\widetilde{X}_{i-1})).$
The first factor satisfies $\nu(\cdot-1)\geq 1$: indeed,
$E_{i,p}((pX_{0},0,\ldots,0))=E_{i,p}(pX_{0})=\exp(T)$
with
$T=\sum_{r=0}^{i-1}(\zeta_{p^{i-r}}-1)\frac{(pX_{0})^{p^{r}}}{p^{r}}=\sum_{r=0}^{i-1}(\zeta_{p^{i-r}}-1)\,p^{p^{r}-r}\,X_{0}^{p^{r}}.$
Each coefficient of $T$ has valuation at least $1$, hence $T\in p\,\mathbb{Z}_{(p)}[\zeta_{p^{i}}]\llbracket X_{0}\rrbracket$ and so $\exp(T)\equiv 1\pmod{p}$. By the induction hypothesis applied to $E_{i-1,p}$ and the congruences $\widetilde{X}_{j}\equiv X_{j-1}^{p}\pmod{p}$, we obtain
$\nu\Big(E_{i-1,p}((\widetilde{X}_{1},\ldots,\widetilde{X}_{i-1}))-1\Big)\geq\frac{1}{p^{i-2}(p-1)},$
and this bound is sharp except in the exceptional case $(p,i)=(2,2)$. Consequently,
$\nu\big(E_{i,p}(p\underline{X})-1\big)=\frac{1}{p^{i-2}(p-1)}\qquad\text{for }(p,i)\neq(2,2).$
Step 4: deduce the bound for $A$ (hence for $E_{i,p}(\underline{X})-1$). For $(p,i)\neq(2,2)$, we have $\nu(E_{i,p}(p\underline{X})-1)<1$, hence in particular $\nu(A)<1$. Also $\nu\!\left(\binom{p}{j}\right)=1$ for $1\leq j\leq p-1$.
We first note that in fact $\nu(A)<\frac{1}{p-1}$: if $\nu(A)\geq\frac{1}{p-1}$, then
$\nu(A^{p})\geq\frac{p}{p-1}>1\qquad\text{and}\qquad\nu\!\left(\binom{p}{j}A^{j}\right)=1+j\nu(A)\geq 1+\frac{1}{p-1}=\frac{p}{p-1}>1,$
so every term in $A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j}$ has valuation >1>1, contradicting $\nu(E_{i,p}(p\underline{X})-1)<1$.
Thus $\nu(A)<\frac{1}{p-1}$, and consequently
$\nu\!\left(\binom{p}{j}A^{j}\right)=1+j\nu(A)\geq 1+\nu(A).$
Since $(p-1)\nu(A)<1$, we have $p\nu(A)<1+\nu(A)$, so the minimal valuation in $A^{p}+\sum_{j=1}^{p-1}\binom{p}{j}A^{j}$ is $p\nu(A)$. Therefore
$\nu\big(E_{i,p}(p\underline{X})-1\big)=p\nu(A),\qquad\text{hence}\qquad\nu(A)=\frac{1}{p^{i-1}(p-1)}.$
This proves that every nonconstant term of $E_{i,p}((X_{0},\ldots,X_{i-1}))$ has valuation $\geq\frac{1}{p^{i-1}(p-1)}$, and the bound is sharp.
Finally, a direct computation in the exceptional case $(p,i)=(2,2)$ shows $\nu\big(E_{2,p}((X_{0},X_{1}))-1\big)=\frac{1}{2}$ (with $p=2$). ∎
@@
Suppose $f(X)=\sum_{j\geq 0}a_{j}X^{j}\in R\llbracket X\rrbracket$, with $a_{0}\in R^{\times}$, $\nu(a_{j})>0$ for $j>0$ and the valuations bounded below by a number $B>0$. Suppose moreover that $f(X)^{p}$ has a finite number of terms whose coefficients have valuations strictly smaller than p/(p1)p/(p-1). Then $f(X)$ itself has only finitely many terms whose coefficients have valuations strictly smaller than 1/(p1)1/(p-1).
@@
Let $u:=a_{0}^{-1}f$. Since $a_{0}\in R^{\times}$, multiplying by the unit $a_{0}^{-p}$ does not change coefficient valuations, so $f^{p}$ has a finite number of terms whose coefficients have valuation $\tfrac{p}{p-1}$ if and only if $u^{p}$ does. Thus we may assume $a_{0}=1$.